### v1

# Violation of local reciprocity in charge–orbital interconversion

Abstract
We demonstrate a violation of local reciprocity in the interconversion between charge and orbital currents. By investigating orbital torque and orbital pumping in W/Ni bilayers, we show that the charge-orbital interconversion in the bulk of the W layer exhibits opposite signs in the direct and inverse processes—the direct and inverse orbital Hall effects being positive and negative, respectively. This finding provides direct evidence of local non-reciprocity in the charge-orbital interconversion, in agreement with a theoretical prediction. These results highlight the unique characteristics of charge-orbital coupled transport and offer fundamental insights into the mechanisms underlying orbital-current-driven phenomena.
Since the discovery of the spin Hall effect (SHE), the interconversion between charge and spin currents through spin-orbit coupling has played a crucial role in the development of spintronics RevModPhys.87.1213 . The direct process of the SHE converts a charge current into a spin current, enabling the electric control of magnetization through spin-orbit torques AndoPRL ; liu2012spin . This process has attracted extensive attention for its topological and quantum mechanical nature, as well as its potential for spintronic applications, such as non-volatile memories and neuromorphic computing RevModPhys.91.035004 . The inverse process of the SHE, the conversion from a spin current into a charge current, offers a way for the electric detection of spin currents saitoh2006conversion ; valenzuela2006direct ; KimuraPRL . This mechanism has lead to the discovery of various spintronic phenomena, including spin pumping—a process in which magnetization dynamics generates a spin current RevModPhys.87.1213 .
While spintronics has primarily focused on the generation and conversion of spin currents, recent advances have revealed that charge currents can also drive the nonequilibrium dynamics of orbital angular momentum (OAM). A key phenomenon underlying OAM dynamics is the direct orbital Hall effect (OHE)—the orbital counterpart of the direct SHE—in which a charge current generates an orbital current, i.e., a flow of electrons carrying nonzero OAM (see Fig. 1)PhysRevLett.95.066601 ; PhysRevLett.102.016601 ; PhysRevLett.121.086602 ; PhysRevB.102.035409 ; PhysRevLett.126.056601 . Theoretical and experimental progress on the OHE has led to the emergence of orbitronics, which aims to utilize the interplay between charge and orbital currents in solid-state devices go2021orbitronics ; KIM2022169974 .
Figure 1: Schematic illustration of orbital torque induced by the direct OHE (left) and the inverse OHE induced by orbital pumping (right) in W/Ni. In the direct OHE, $j_{\mathrm{C}}$ represents the applied charge current, while $j_{\mathrm{L}}^{\mathrm{bulk}}$ and $j_{\mathrm{L}}^{\mathrm{surface}}$ denote the orbital currents generated by the bulk and surface contributions, respectively. The red and blue colors represent the positive and negative orbital polarizations, respectively. In the direct OHE, $j_{\mathrm{L}}^{\mathrm{bulk}}$ and $j_{\mathrm{L}}^{\mathrm{surface}}$ carry opposite orbital angular momentum because the bulk and surface orbital Hall conductivities have opposite signs in W. In the inverse OHE, $j_{\mathrm{L}}$ is the injected orbital current, while $j_{\mathrm{C}}^{\mathrm{bulk}}$ and $j_{\mathrm{C}}^{\mathrm{surface}}$ are the charge currents generated by the bulk and surface contributions, respectively. These charge currents, $j_{\mathrm{C}}^{\mathrm{bulk}}$ and $j_{\mathrm{C}}^{\mathrm{surface}}$, flow in opposite directions. $M$ denotes the magnetization.
[…]
$\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)\neq\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z).$ (1)
This prediction indicates that while the direct and inverse OHEs are globally reciprocal, their local reciprocity is violated. This is in stark contrast to the SHE, which has been predicted to be reciprocal even locally go2024local . Such a distinction is of fundamental importance, as it provides a key experimental signature for differentiating between spin and orbital currents. However, despite recent observations of the direct and inverse OHEs Cr-orbital ; Ta-orbital ; PhysRevResearch.4.033037 ; hayashi2022observation ; choi2021observation ; hayashi2024observation ; hayashi2022observation ; hayashi2023orbital ; choi2021observation ; PhysRevB.107.134423 ; xu2024orbitronics ; hayashi2024observation ; seifert2023time ; gao2024control ; 4c06056 , the nature of their reciprocal relationship has remained unclear. In nanoscale spintronic and orbitronic devices, the measured quantities do not exactly correspond to macroscopic averages. In such devices, local reciprocity may be investigated by carefully measuring spin and orbital responses across devices with systematically varied thicknesses. This approach could offer a pathway to exploring coupled transport phenomena beyond the framework of Onsager’s reciprocity.
In this Letter, we report the observation of a violation of local reciprocity in the charge-orbital interconversion. We investigate the direct and inverse OHEs in W using orbital torque and orbital pumping in W/Ni bilayers (see Fig. 1PhysRevResearch.2.013177 ; hayashi2024observation . In this system, the charge-to-spin/orbital conversion can be investigated by measuring current-induced spin and orbital torques on the Ni magnetization hayashi2022observation . The reciprocal process, the spin/orbital-to-charge conversion is induced by ferromagnetic resonance (FMR). The FMR drives the orbital pumping, as well as the spin pumping, due to the strong spin-orbit correlation of Ni hayashi2024observation ; arXiv:2309.14817 ; han2023theory ; PhysRevApplied.19.014069 ; el2023observation , injecting spin and orbital currents into the W layer. These currents are subsequently converted into charge currents via the inverse SHE and OHE, respectively. Through systematic measurements of the orbital torque and orbital pumping, we demonstrate that the OHE in the W bulk exhibits opposite signs in the direct and inverse processes. This asymmetry aligns with a theoretical prediction of the strong local non-reciprocity in the OHEs of W go2024local , revealing the unique characteristics of charge-orbital coupled transport.
To investigate the charge-to-orbital conversion in W, we measured the current-induced torque in the W/Ni bilayers using spin-torque ferromagnetic resonance (ST-FMR). The samples were fabricated by magnetron sputtering with the structure of SiO2(4 nm)/W($t_{\mathrm{W}}$)/Ni($t_{\mathrm{Ni}}$)/SiO2-substrate , where the numbers in parentheses indicate the layer thicknesses. The measurement and analysis procedures for the ST-FMR follow those in our previous work hayashi2022observation . In Fig. 2(a), we show the W-layer thickness $t_{\mathrm{W}}$ dependence of the damping-like torque efficiency, $\xi_{\text{DL}}^{E}=(2e/\hbar)\mu_{0}M_{\text{s}}t_{\text{Ni}}H_{\text{DL}}/E$, determined by the ST-FMR for the W/Ni bilayers, where $M_{\mathrm{s}}$ is the saturation magnetization, $H_{\mathrm{DL}}$ is the damping-like effective field, and $E$ is the applied electric field.
Figure 2: The damping-like torque efficiency $\xi_{\mathrm{DL}}^{E}$ determined by the ST-FMR for the W/Ni bilayers. (a) Dependence of $\xi_{\mathrm{DL}}^{E}$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer. (b) Dependence of $\xi_{\mathrm{DL}}^{E}$ on the Ni thickness $t_{\rm Ni}$ in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
Figure 2(a) shows that $\xi_{\text{DL}}^{E}$ changes sign from negative to positive as $t_{\mathrm{W}}$ increases, indicating a crossover in the dominant mechanism responsible for the observed torque. This result is consistent with our previous observation hayashi2022observation . The negative sign of $\xi_{\mathrm{DL}}^{E}$ aligns with that of the SHE in the W layer PhysRevMaterials.6.095001 , suggesting the direct SHE is the primary mechanism driving this torque. We note that the increase in $\xi_{\mathrm{DL}}^{E}$ with $t_{\mathrm{W}}$ indicates that the torque with a positive sign becomes pronounced as $t_{\mathrm{W}}$ increases, suggesting that this torque originates from the charge current flowing in the bulk of the W layer. This $t_{\mathrm{W}}$-dependent behavior is consistent with what is expected from the bulk direct OHE in the W layer.
The significant contribution of the orbital torque to $\xi_{\mathrm{DL}}^{E}$ in the W/Ni bilayer is supported by a long-range nature of $\xi_{\mathrm{DL}}^{E}$. A key difference between spin and orbital phenomena lies in the characteristic length scales associated with the thickness of the ferromagnet (FM) PhysRevLett.130.246701 . It has been demonstrated that spin torque efficiency is nearly independent of the FM-layer thickness due to the short-range nature of spin transport in FMs, whereas orbital torque efficiency increases with FM-layer thickness because of its long-range nature PhysRevLett.130.246701 ; hayashi2022observation ; PhysRevLett.128.067201 ; PhysRevB.107.134423 ; PhysRevResearch.4.033037 ; PhysRevB.105.104434 ; PhysRevResearch.5.023054 ; gao2024control . This difference in the length scales arises because spin currents decay rapidly—typically within less than 1 nm–due to spin dephasing in FMs, whereas orbital currents can propagate over much longer distances through momentum-space hotspots, which suppress oscillations of the OAM PhysRevLett.130.246701 . As shown in Fig. 2(b), $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, even though the $t_{\mathrm{Ni}}$ range is much larger than the spin dephasing length. This result suggests that $\xi_{\mathrm{DL}}^{E}$ with the positive sign is dominated by the orbital torque induced by the bulk direct OHE. Here, Figure 2(b) also shows that in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ and changes sign. This result can be interpreted as arising from a competition between spin and orbital torques. While the spin torque due to the direct SHE with the negative sign provides the dominant contribution to $\xi_{\mathrm{DL}}^{E}$ in the W(3 nm)/Ni(5 nm) bilayer, this contribution remains nearly independent of $t_{\mathrm{Ni}}$. In contrast, the orbital torque with the positive sign due to the direct OHE increases with $t_{\mathrm{Ni}}$, leading to the sign reversal of $\xi_{\mathrm{DL}}^{E}$.
The W layer in the W/Ni bilayer predominantly consists of $\alpha$-W, which is confirmed by the dependence of the W-layer resistivity $\rho_{\rm W}$ on $t_{\rm W}$, shown in Fig. 3(a). Figure 3(a) shows that $\rho_{\rm W}$ follows $\rho_{\rm W}(t_{\rm W})=at_{\rm W}^{-1}+\rho_{\rm W}^{\rm bulk}$, where $at_{\rm W}^{-1}$ accounts for the resistivity contribution from surface scattering, and $\rho_{\rm W}^{\rm bulk}$ is the bulk-limit resistivity. The extracted value of $\rho_{\rm W}^{\rm bulk}=17~{}\mu\Omega\text{cm}$ is consistent with the resistivity of low-resistivity $\alpha$-W hayashi2022observation . These results indicate that the sign of the bulk direct orbital Hall conductivity $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}$ is positive in $\alpha$-W. This observation is consistent with theoretical predictions PhysRevLett.102.016601 ; PhysRevMaterials.6.095001 ; go2024local .
In the following, we explore the orbital-to-charge conversion in the $\alpha$-W layer by investigating the inverse process of the orbital torque generation—orbital pumping. For the pumping measurements, the W/Ni film was patterned into rectangular strips with a width of $8$ $\mu$m and a length of $180$ $\mu$m using photolithography and Ar-ion milling. The device was positioned between the ground and signal lines of a Ti/Au coplanar waveguide (CPW), with a gap of 8 $\mu$m separating the CPW from the device (see Fig. 3(b)). During the measurements, a radio frequency (RF) current $I_{\mathrm{RF}}$ with a frequency $f$ and a power $P_{\mathrm{in}}$ was applied to the CPW, along with an in-plane external magnetic field $H$ at an angle $\theta$ relative to the longitudinal direction of the device. The RF current generates an out-of-plane microwave magnetic field, driving FMR in the W/Ni bilayer. We measured the direct-current (DC) voltage $V_{\rm DC}$ induced by the FMR across electrodes attached to the W/Ni bilayer.
Figure 3: Charge-current signals induced by spin/orbital pumping. (a) Dependence of the W-layer resistivity $\rho_{\mathrm{W}}$ on $t_{\rm W}$ in the Ni(5 nm)/W($t_{\mathrm{W}}$) bilayer. The solid circles are the experimental data, and the solid cure is the fitting result. (b) Schematic illustration of the experimental setup used to measure the spin/orbital pumping. ${\bf H}$ and ${\bf I}_{\mathrm{RF}}$ denote the external magnetic field and the RF current, respectively. $\theta$ is defined as the in-plane angle between ${\bf H}$ and ${\bf I}_{\mathrm{RF}}$. $H$ dependence of $I_{\rm DC}$ for the (c) W(0.4 nm)/Ni(5 nm) and (d) W(6 nm)/Ni(5 nm) bilayers at $\theta=90^{\circ}$, with $f=10$ GHz and $P_{\rm in}=200$ mW. The solid circles represent the experimental data, and the black solid curves show the fitting results, which are composed of the sum of symmetric (red curve) and antisymmetric (blue curve) components. $\theta$ dependence of $I_{\rm{sym}}$ for the (e) W(0.4 nm)/Ni(5 nm) and (f) W(6 nm)/Ni(5 nm) bilayers. The solid circles represent the experimental data, and the black solid curves show the fitting results using Eq. (2).
In Figs. 3(c) and 3(d), we show the dependence of the charge current $I_{\rm DC}=V_{\rm DC}/R$ on $H$ for the W(0.4 nm)/Ni and W(6 nm)/Ni bilayers, respectively, where $R$ denotes the device resistance. By fitting the measured signal using $I_{\rm{DC}}=I_{\rm{sym}}{(\mu_{0}W)^{2}}/[{(\mu_{0}H-\mu_{0}H_{\rm{res}})^{2}+(\mu_{0}W)^{2}}]+I_{\rm{asym}}{(\mu_{0}W)(\mu_{0}H-\mu_{0}H_{\rm{res}})}/[{(\mu_{0}H-\mu_{0}H_{\rm{res}})^{2}+(\mu_{0}W)^{2}}]$, we extract the symmetric component $I_{\rm{sym}}$, where $H_{\rm{res}}$ is the FMR field and $W$ is the FMR linewidth. Here, the spin and orbital pumping generates $I_{\rm{sym}}$, while spin rectification effects, including the anisotropic magnetoresistance (AMR), the planar Hall effect (PHE), and the anomalous Hall effect (AHE) in the Ni layer, can contribute to both the symmetric $I_{\rm{sym}}$ and antisymmetric $I_{\rm{asym}}$ components.
To extract the pumping signal $I_{\mathrm{pump}}$, we measured $I_{\mathrm{DC}}$ by varying the in-plane magnetic field angle $\theta$, as shown in Figs. 3(e) and 3(f). The results are consistent with HARDER20161
$I_{\mathrm{sym}}=I_{\mathrm{pump}}\sin\theta+I_{\mathrm{AMR}}\sin 2\theta+I_{\mathrm{PHE}}\cos 2\theta+I_{\mathrm{AHE}},$ (2)
where $I_{\mathrm{AHR}}$, $I_{\mathrm{PHE}}$, and $I_{\mathrm{AHE}}$ are the spin rectification signals due to the AMR, PHE, and AHE, respectively. Figures 3(e) and 3(f) demonstrate that the sizable $I_{\mathrm{pump}}$ signals are generated by the spin and/or orbital pumping in the W/Ni bilayers. Using the extracted value of $I_{\mathrm{pump}}$, we determine the pumping-induced charge-current generation efficiency $\zeta$, defined as supplementary
$\zeta=I_{\mathrm{pump}}\frac{\gamma\mu_{0}H_{\mathrm{res}}\left({\mu_{0}W}\right)^{2}}{2\pi ewf^{2}}\left[1+\left(\frac{2\pi f}{\gamma\mu_{0}H_{\mathrm{res}}}\right)^{2}\right]^{2},$ (3)
where $\gamma$ is the gyromagnetic ratio, $e$ is the elementary charge, and $w$ is the width of the device. In the model of the spin pumping and the inverse SHE, $\zeta$ represents an efficiency parameter that characterizes the strength of the spin pumping and the inverse SHE as $\zeta=g_{\mathrm{eff}}^{\uparrow\downarrow}\theta_{\mathrm{H}}\lambda\tanh\left[{t_{\mathrm{W}}}/(2\lambda)\right](\mu_{0}h)^{2}$, which varies with the W thickness over the spin diffusion length but should be independent of the Ni thickness. Here, $g_{\mathrm{eff}}^{\uparrow\downarrow}$ is the effective spin-mixing conductance, $\theta_{\mathrm{H}}$ is the spin Hall angle, and $\lambda$ is the spin diffusion length. $h$ is the amplitude of the microwave magnetic field, which is consistent across all devices used in this study.
Figure 4: The pumping-induced charge-current generation efficiency $\zeta$ in the W/Ni bilayers. (a) Dependence of $\zeta$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer. (b) Dependence of $\zeta$ on the Ni thickness $t_{\rm Ni}$ in the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
In Fig. 4(a), we show the dependence of the pumping efficiency $\zeta$ on the W-layer thickness $t_{\mathrm{W}}$. Notably, $\zeta$ remains negative in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer even for $t_{\mathrm{W}}>6$ nm, where the orbital torque efficiency $\xi_{\mathrm{DL}}^{E}$ becomes positive (see Fig. 2(a)). The negative sign of $\zeta$ corresponds to the negative sign of the charge-current generation mechanism in the W layer. This implies that $\zeta$ at $t_{\mathrm{W}}>6$ nm is dominated by the inverse SHE, which is known to be negative in W. Nevertheless, as shown in Fig. 4(b), $\zeta$ exhibits a strong dependence on the Ni-layer thickness $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer. This behavior demonstrates that $\zeta$ cannot be solely attributed to the spin pumping, which should be independent of $t_{\mathrm{Ni}}$, just as the spin torque—the reciprocal process of the spin pumping—is also independent of $t_{\mathrm{Ni}}$. We further confirm that the pumping signal is negligible in Ni single-layer films within this $t_{\mathrm{Ni}}$ range supplementary .
The observed long-range dependence of $\zeta$ on $t_{\mathrm{Ni}}$ is characteristic of the reciprocal process of the long-range orbital torque, indicating that the orbital pumping contributes to the charge current in the W(10 nm)/Ni bilayer. Notably, the decrease in $\zeta$ with $t_{\mathrm{Ni}}$ indicates that the orbital-to-charge conversion has a negative sign. In addition to this negative orbital-to-charge conversion, we also observe a positive orbital-to-charge conversion in the W/Ni bilayer. As shown in Fig. 4(a), $\zeta$ exhibits a positive peak around $t_{\mathrm{W}}=0.5$ nm. This positive component increases with $t_{\mathrm{Ni}}$ as shown in Fig. 4(b), indicating that the signal arises from the orbital-to-charge conversion with a positive sign induced by the orbital pumping. These results demonstrate that two distinct orbital-to-charge conversion mechanisms with opposite signs coexist in the W/Ni bilayer: a positive contribution for $t_{\mathrm{W}}<1$ nm and a negative contribution for $t_{\mathrm{W}}>1$ nm.
The orbital-to-charge conversion with the negative sign can be attributed to the bulk inverse OHE in the W layer, demonstrating a violation of local reciprocity between the direct and inverse OHEs. The nature of the orbital-to-charge conversion can differ between the bulk and interfaces, accounting for the two orbital-to-charge conversion mechanisms with different signs observed in the W/Ni bilayer. Since the interface contribution originates from the anisotropic crystal potential, which differs from the bulk crystal potential go2024local , we assume that the interface contribution is particularly significant at the SiO2/W interface rather than at the metallic W/Ni interface. This interface effect should be especially strong at small $t_{\mathrm{W}}$, suggesting that the charge-current generation with the positive sign at $t_{\mathrm{W}}<1$ nm can be attributed to the orbital-to-charge conversion at the SiO2/W interface. As $t_{\mathrm{W}}$ increases, this surface contribution is suppressed, while bulk effects, including the bulk inverse SHE and OHE, become more pronounced. This indicates that the bulk inverse OHE dominates the observed orbital-to-charge conversion with the negative sign at larger $t_{\mathrm{W}}$. These results demonstrate that while the sign of the bulk direct OHE is positive, as evidenced by the torque measurements, the pumping measurements show that the sign of the bulk inverse OHE is negative. The observation of opposite signs for the bulk direct and inverse OHEs demonstrates that the local reciprocity is violated in the charge-orbital interconversion in the $\alpha$-W.
Table 1: Decomposition of the bulk and surface contributions to the direct and inverse OHEs in W go2024local . The bulk and surface contributions to the direct(inverse) orbital Hall conductivity are represented as $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),surface}}$, respectively, along with the spin Hall conductivity components represented as $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),surface}}$.
Orbital/Spin Conversion Bulk Surface
OHE Direct $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}<0$
Inverse $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{OH}}^{\mathrm{inv,surface}}>0$
SHE Direct $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}\simeq 0$
Inverse $\sigma_{\mathrm{SH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{inv,surface}}\simeq 0$
Our observations are consistent with the theoretical prediction of a violation of local reciprocity between the direct and inverse OHEs in $\alpha$-W go2024local . In the theory, the non-reciprocity in the charge-orbital interconversion arises from the non-conservation of the OAM due to its strong interaction with the lattice via crystal fields. This interaction is accounted for by defining the orbital current as the sum of the conventional orbital current and the torque dipole, the latter being associated with the interaction between the OAM and the crystal lattice. It has been shown that the torque dipole is particularly significant at interfaces, where the anisotropic crystal potential efficiently mediates angular momentum exchange between the electrons and the lattice, strongly influencing the local profile of the charge-orbital interconversion in the W layer. As a result, the local charge currents induced by the orbital current deviate from the expectation of strict reciprocity. Notably, even the signs of the local responses differ between the direct and inverse OHEs, as summarized in Table I. The positive and negative signs of the bulk direct and inverse orbital Hall conductivities, $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ and $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$, evidenced by the torque and pumping results, are consistent with this prediction. Importantly, these opposite signs do not violate Onsager’s reciprocity relation, as the direct and inverse responses are reciprocal when summed over the entire film thickness go2024local . These behaviors contrasts sharply with the SHE in $\alpha$-W, where the direct and inverse SHEs are largely reciprocal even at a local level go2024local . This indicates that the observed violation of the local reciprocity between the direct and inverse OHEs highlights the distinct nature of the charge-orbital interconversion compared to the charge-spin interconversion. Table I also suggests that the surface direct OHE with a negative sign may contribute to the negative torque for $t_{\mathrm{W}}<4$ nm.
In summary, we have experimentally demonstrated the violation of local reciprocity in the charge-orbital interconversion in the W/Ni bilayers. By measuring the orbital torque and orbital pumping, we found that the sign of the bulk inverse OHE is opposite to that of the bulk direct OHE. This result confirms that the charge-orbital interconversion is locally non-reciprocal, in sharp contrast to the global reciprocity governed by Onsager’s reciprocal relations. The observation of the local reciprocity violation in the direct and inverse OHEs highlights the unique characteristics of charge-orbital coupled transport, providing essential insights for a deeper understanding of orbitronic phenomena.
Acknowledgements.
This work was supported by JSPS KAKENHI (Grant Number: 22H04964, 23K19037), Spintronics Research Network of Japan (Spin-RNJ), and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (Grant Number: JPJ011438). Y.M. and D.G. acknowledge support from the EIC Pathfinder OPEN grant 101129641 “OBELIX”.